40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 21 18:37:16 2009, Jenne, Update, General, Transfer function of Mode Cleaner Stacks TF_osems_guralps.png
    Reply  Tue Aug 25 00:35:04 2009, rana, Update, General, Transfer function of Mode Cleaner Stacks Untitled.png
Message ID: 1938     Entry time: Tue Aug 25 00:35:04 2009     In reply to: 1935
Author: rana 
Type: Update 
Category: General 
Subject: Transfer function of Mode Cleaner Stacks 

Looks like all of the accelerometers and seismometers have been disconnected since early AM last Monday when Clara disconnected them for her sensor noise measurement.

Attachment 1: Untitled.png  24 kB  | Hide | Hide all
Untitled.png
ELOG V3.1.3-